OpenStudy (anonymous):

Justin wants to use 376 ft of fencing to fence off the greatest possible rectangular area for a garden. What dimensions should he use? What will be the area of the garden? A. 89 x 99; 8811 ft B. 92 x 96; 8832 ft C. 94 x 94; 8836 ft D. 93 x 95; 8835 ft please don't just provide the answer I would like to know how, will fan and medal

OpenStudy (anonymous):

http://www.cymath.com/

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

The perimeter of the rectangle is P = 2*(L+W). Plug in P = 376 and isolate L P = 2*(L+W) 376 = 2*(L+W) 376/2 = 2*(L+W)/2 188 = L+W L+W = 188 L = 188-W now plug this into A = L*W A = L*W A = (188-W)*W A = 188W - W^2 A = -W^2 + 188W If you replace A with y, and replace W with x, you will get this equation y = -x^2 + 188x finding the vertex of that equation leads you to the max area (and the dimensions needed to get the max area)